OpenStudy (anonymous):

Need help applying compound interest formula:

OpenStudy (anonymous):

to what?

OpenStudy (anonymous):

H! So I have to apply the formula to a pretend product. The product costs 988.00 and has a 29.99% interest rate. I need to find how much my balence will be at the end of the deferment period.

OpenStudy (anonymous):

B = ( P (1 + i)^nt ) So I will use this formula.. B= unknown P=988 i= rate/number of compounding periods per year = .2999/12 (the interest is compounded monthly) t= 2.75 (I have a "2 year 9 month deferment promo plan")

OpenStudy (anonymous):

Here is where I am B=988 (1+.2999)^33 B= 5672638 This doesn't seem possible, what am I doing wrong?

terenzreignz (terenzreignz):

Are you sure it's not 0.2999% a 29.99 % rate of interest is highly unrealistic XD

OpenStudy (anonymous):

"or new accounts: Purchase APR is 29.99%; "

terenzreignz (terenzreignz):

I see... hang on. Actual formula if memory serves is this: \[\Large B = P \left(1+\frac{r}{n}\right)^{nt}\] But you forgot to divide 0.2999 by 12...

OpenStudy (anonymous):

ahh, jeez! Thanks! So the balence should be 2231.15

terenzreignz (terenzreignz):

Does that seem more realistic to you? :D